Can you help me please.
f(x)=5x+4; g(x)=7x-1 and solve for (f-g)(x)
Thank you

Answer by fractalier(6550)

(f-g)(x) = f(x) - g(x) = 5x+4 - (7x-1) = 5x + 4 - 7x + 1 = -2x + 5